Circular No. 8014

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hw IN UGC 52 M. Schwartz and W. Li report the LOTOSS (cf. IAUC 7906) discovery

of an apparent supernova on unfiltered images taken with the 0.8-m Tenagra II automated telescope on Nov. 9.2 (mag ∼ 16.8) and 10.2 UT (mag ∼ 16.6). SN 2002hw is located at α = 0^{h}06^{m}49^{s}.06, δ = +8^{o}37′48′′.5 (equinox 2000.0), which is 6′′.7 west and 6′′.0 north of the nucleus of UGC 52. A KAIT image from Nov. 2.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVAE 2002gz AND 2002ho A. V. Filippenko, R. Chornock, and S. Jha, University of California,

Berkeley; and B. Barris, University of Hawaii, write: “Inspection of CCD spectra (range 400–1000 nm), obtained on Nov. 9 UT with the Keck II 10-m telescope (+ ESI), shows that SN 2002gz (IAUC 7996) is indeed a supernova, but its spectral classification is uncertain at this time. It most closely resembles type II, with a blue continuum and unusually broad undulations at the approximate positions expected of hydrogen Balmer lines and O I 777.4-nm. Moreover, there may be He I absorption superposed on the Hα emission, so perhaps its spectral type is actually IIb. The redshift of the host galaxy is 0.085, based on narrow emission lines from adjacent H II regions. SN 2002ho (IAUC 8011) is of type Ic, within a few weeks past maximum brightness; the Ca II near-infrared triplet exhibits an especially prominent and broad P-Cyg profile.”

POSSIBLE NOVA IN CEPHEUS Seiichi Yoshida, Ibaraki, Japan, reports (together with J. Greaves,

Northampton, East Midlands, U.K.; and K. Kadota, Ageo, Saitama, Japan) his discovery, in the course of the Multitudinous Image-based Sky-survey and Accumulative Observations Project, of a new variable star on unfiltered CCD images taken with a 500-mm camera lens by N. Ohkura, Okayama, Japan, on 2001 Sept. 19.56 (at mag 12.7) and Nov. 10.45 UT (mag 14.3). The new object is located at α = 22^{h}58^{m}09^{s}.12, δ = +66^{o}21′12′′.4 (equinox 2000.0). Nothing appears at this position on Palomar Sky Survey plates or Digitized Sky Survey images. Unfiltered CCD images taken by T. Muramatsu (Fukushima, Japan) on 2002 Oct. 25.51 with a 0.65-m reflector show this object at mag ∼ 20. Images taken by others (including A. Takao and M. Collins) show nothing to mag 12.3 on 2001 Sept. 16 and 17 and to mag 13.0 on 2001 June 28 and Oct. 8.

2002 November 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green